OpenStudy (anonymous):

find a value for K so that if f is continuous at x=3 f(x)=...

OpenStudy (anonymous):

\[f(x)= \left\{ \sqrt{5x+1} - \sqrt{x+13} , \ for x \neq 3\right\} and \left\{ k , \, =3\right\}\]

OpenStudy (anonymous):

I'm confused at what the question is asking...

OpenStudy (anonymous):

is the second part suppose to be a point?

OpenStudy (anonymous):

yes k is the point that i have to find a value for to make the function continuous, {k, for when x=3}

OpenStudy (anonymous):

do you mean when the function = 3?

OpenStudy (anonymous):

I'm a little confused.

OpenStudy (anonymous):

your point is then (3,k)?

OpenStudy (anonymous):

i believe so

OpenStudy (anonymous):

ok. then x=3 and y=k

OpenStudy (anonymous):

plug in 3 and solve for y